# Koopman Autoencoders with Continuous-Time Latent Dynamics for Fluid Dynamics Forecasting

### v1

@@ 6.1 Experimental Setup
Standard training of ACDM utilizes non-overlapping sequences to form training batches. This approach discards a significant portion of valid temporal transitions. In contrast, we adopt an exhaustive sliding-window strategy, where we generate training batches by sliding the window forward by a single time step. This implies that if a sequence begins at $t_{i}$, the next valid training sequence begins at $t_{i+1}$. This strategy significantly increases the effective dataset size and forces the encoder to learn a more robust, temporally consistent manifold.
@@
The probabilistic nature of the diffusion baseline constrains its objective function to a simple MSE objective derived from the Gaussian diffusion assumption. Our deterministic framework allows for a composite loss function (as defined in (16)) that incorporates physics-informed constraints, leading to sharper structural predictions.
@@ 6.2 Results
We structure our evaluation into four main parts: a quantitative analysis comparing our method against baselines, a validation of the learned integrator against the analytical matrix exponential, an assessment of temporal consistency across varying resolutions, and a final study on computational efficiency.
Quantitative Analysis.
Table 1 illustrates the results of five distinct models evaluated on the full Incompressible Flow dataset. The comparison reveals that the three models trained exclusively on the fully developed regime exhibit significantly higher MSE, primarily due to their lack of exposure to the initial transient dynamics. Most notably, our Full Continuous KAE outperforms Full retrained ACDM in both $Inc_{low}$ and $Inc_{high}$.
Table 3 details the quantitative performance evaluated on fully developed Incompressible Flow and full dataset of Transonic Flow. Our continuous KAE achieves predictive accuracy highly competitive with ACDM baseline, yielding MSE and LSiM scores effectively comparable with ACDM. Figure 2 further provides a visual comparison of rollout predictions on $Inc_{high}$ and $Tra_{ext}$, where $t$ denotes the number of autoregressive steps (corresponding to physical time $t\cdot\Delta t$). Notably, the chaotic nature of the transonic flow proves challenging for long-horizon forecasting, with both ACDM and Continuous KAE starting to exhibit visible error accumulation and divergence from the ground truth at around step $t=45$.
@@
We compare the trajectories generated by the numerical RK4 integrator against the analytical matrix exponential solution derived in Section 3.2 at inference time. Results are shown in Figure 3 and the final two rows of Table 3. While RK4 performs iterative rollouts with a fixed step $\Delta t=0.1s$, the analytical formulation allows us to directly compute $z_{\tau}=\exp(\mathbf{K}\tau)z_{0}$ for any target time $\tau$. In Figure 3, $t$ denotes the number of rollout steps, corresponding to a physical time $\tau=t\cdot\Delta t$. Both Figure 3 and Table 3 show perfect alignment with two methods. This allows for highly efficient long-horizon forecasting, because to predict a frame far into the future, one can simply utilize the matrix exponential form to compute the state in a single step, bypassing the computational overhead of iterative integration.
@@
A key advantage of our continuous-time KAE is its ability of zero-shot temporal generalization. Despite being trained on a fixed timestep $\Delta t=0.1s$, the model supports arbitrary integration step sizes during inference. We validate this capability in Figure 4, which displays rollouts generated using three distinct integration step sizes, $\Delta t\in\{0.05s,0.1s,0.2s\}$. For each case, the number of rollout steps is adjusted to align with the same set of physical time instants (up to $t=2.0s$). The first row also shows the results produced using the analytical exponential form. The perfect alignment confirms that the learned dynamics are robust to discretization changes.
@@
Action / Metric ACDM KAE
Encode (ms $\times 10^{-1}$) $2.44$
Decode (ms $\times 10^{-2}$) $3.56$
240-step rollout (ms $\times 10^{-1}$) $341.42$ $1.21$
Latent avg./step (ms $\times 10^{-4}$) $5.06$
Total avg./step (ms $\times 10^{-3}$) $142.26$ $1.67$
@@
$\text{ACDM}_{ncn}$ $\mathbf{0.9\pm 0.8}$ $6.6\pm 2.7$ $5.7\pm 2.7$ $3.6\pm 1.2$ $4.1\pm 1.9$ $1.9\pm 0.6$ $2.8\pm 1.3$ $1.7\pm 0.4$ $22.2\pm 3.8$ $8.2\pm 3.0$
ACDM $1.7\pm 2.2$ $6.9\pm 5.6$ $\mathbf{0.8\pm 0.4}$ $\mathbf{1.0\pm 0.3}$ $\mathbf{2.3\pm 1.4}$ $\mathbf{1.3\pm 0.3}$ $\mathbf{2.7\pm 2.1}$ $\mathbf{1.3\pm 0.6}$ $22.3\pm 4.6$ $\mathbf{3.8\pm 0.4}$
Continuous KAE $1.3\pm 1.7$ $\mathbf{6.1\pm 4.8}$ $2.9\pm 1.1$ $1.7\pm 0.3$ $2.5\pm 0.8$ $1.8\pm 0.3$ $6.5\pm 1.6$ $2.2\pm 0.4$ $\mathbf{17.0\pm 2.3}$ $4.1\pm 0.3$
Exponential Continuous KAE $1.3\pm 1.7$ $\mathbf{6.1\pm 4.8}$ $2.9\pm 1.1$ $1.7\pm 0.3$ $2.5\pm 0.8$ $1.8\pm 0.3$ $6.5\pm 1.6$ $2.2\pm 0.4$ $\mathbf{17.0\pm 2.3}$ $4.1\pm 0.3$
@@ 7 Discussion
In this work, we presented a continuous-time Koopman Autoencoder that effectively decouples the learning of latent dynamics from the temporal resolution of the training data. Our results on incompressible and transonic flows demonstrate that this approach offers a compelling alternative to state-of-the-art probabilistic methods like ACDM.
Accuracy vs. Efficiency Trade-off. While diffusion models such as ACDM excel at capturing high-frequency texture through stochastic refinement, they incur a significant computational cost during inference due to iterative denoising steps. Our method, by contrast, utilizes a deterministic ODE solver in the latent space. As shown in Table 1, our continuous formulation achieves competitive MSE and LSiM scores, often within the error margins of ACDM, while being two orders of magnitude faster during inference. This efficiency is critical for real-time control and design optimization loops.
@@
Limitations. Despite these successes, we observe that the spectral bias of the $L_{2}$ reconstruction loss can lead to the smoothing of fine-scale turbulent features over very long horizons, a common issue in deterministic autoencoders. Additionally, while the Koopman operator provides interpretability, the assumption of global linearity (even with residual corrections) remains a constraint for highly non-linear phenomena like shock wave propagation in the transonic regime. We also noted that stabilizing the training of the operator requires careful curriculum learning and spectral normalization, indicating a sensitivity to hyperparameter tuning.
@@ 8 Conclusion
We have introduced a Koopman Autoencoder with continuous-time latent dynamics for data-driven fluid dynamics forecasting. By modeling the evolution of flow fields as an ODE in a compressed latent space, our approach combines the theoretical grounding of Koopman operator theory with the flexibility of neural differential equations. We demonstrated that integrating these dynamics with a fourth-order Runge-Kutta scheme allows for accurate, stable, and resolution-invariant predictions across incompressible and transonic flow regimes.
Our extensive benchmarks against conditional diffusion models highlight that continuous-time KAEs can match state-of-the-art accuracy while offering superior computational efficiency and temporal generalization. Future work will focus on scaling this approach to three-dimensional turbulence and incorporating explicit conservation laws into the latent ODE structure to further improve long-term physical fidelity.
@@
This paper presents a method for accelerating the simulation of complex fluid dynamics, which has broad applications in engineering, climate science, and energy sectors. By providing a surrogate model that is orders of magnitude faster than traditional numerical solvers (DNS/LES), this work contributes to more energy-efficient computational workflows, reducing the carbon footprint associated with high-performance computing.
In engineering contexts, such as aerodynamic design or wind energy optimization, faster simulation tools can accelerate the development of efficient technologies. We do not foresee immediate negative ethical consequences or societal risks associated with this specific algorithmic contribution.
@@
Datasets are detailed at https://github.com/tum-pbs/autoreg-pde-diffusion/tree/main . These data sets are available at the training and evaluation resolution of $128\times 64$ (about 146 GB).

### v2

@@ 6.1 Experimental Setup
Standard training of the ACDM baseline utilizes non-overlapping sequences to form training batches. This approach discards a significant portion of valid temporal transitions. In contrast, similar to established data augmentation practices in autoregressive surrogate modeling (Brandstetter et al., 2022; Wu et al., 2022), we adopt an exhaustive sliding-window strategy. By sliding the temporal window forward by a single time step, we generate heavily overlapping batches (i.e., if a sequence begins at $t_{i}$, the next valid training sequence begins at $t_{i+1}$). The sliding-window strategy increases the number of observed temporal transitions, improving supervision of the latent dynamics. Importantly, this does not introduce additional simulation data, but reuses existing trajectories to provide denser temporal coverage. As a result, the model learns a more consistent latent representation while maintaining a fair comparison with baselines.
@@
While diffusion models like ACDM can be modified to include physics-informed guidance during the sampling phase, their standard training objective inherently relies on a probabilistic MSE derived from the Gaussian diffusion assumption. In contrast, our deterministic framework naturally accommodates a direct, composite loss function (as defined in (14)) during the primary training loop. This allows us to strictly enforce physics-informed constraints—such as Sobolev norms and spectral alignment—leading to sharper structural predictions without requiring auxiliary guidance models at inference.
@@ 6.2 Results
We structure our evaluation into five main parts: a quantitative analysis comparing our method against the baselines, a validation of the learned integrator against the analytical matrix exponential, an assessment of temporal consistency across varying resolutions, and a final study on computational efficiency.
We compare against ACDM as a strong generative baseline, alongside autoregressive and Koopman-based models, to evaluate trade-offs between expressivity, stability, and efficiency. Benchmarking against it allows us to explicitly characterize the spectrum of trade-offs in PDE forecasting: the high-frequency synthesis and short-term accuracy of diffusion models versus the temporal flexibility, extreme long-horizon stability, and computational efficiency of continuous-time Koopman theory. Generative models achieve lower short-term error by modeling fine-scale stochastic structure. In contrast, our method produces smoother predictions due to the linear structure of the latent dynamics, which can lead to slightly higher error at short horizons.
We also compare training and inference efficiency across models. While all methods are trained on the same underlying simulation data, their computational costs differ significantly due to their respective optimization objectives and architectures. Diffusion-based models require iterative noise conditioning and sampling, resulting in longer training and inference times. In contrast, our deterministic formulation enables substantially faster convergence and evaluation.
In our experiments, the continuous-time KAE converges in approximately 200 epochs, compared to 1000 epochs for the diffusion baseline, corresponding to wall-clock training times of approximately $24$ hours versus $72$ hours on a single NVIDIA RTX 4090 GPU. We further observe a significant reduction in inference cost, as our model requires only a single forward pass per timestep.
Quantitative Analysis and Phase Alignment.
Tables 1 and 3 detail the quantitative performance across regimes, revealing a clear trade-off between short-term accuracy and long-horizon stability. On the Incompressible Flow dataset (Table 1), the Full Continuous KAE effectively captures the smoother dynamics, outperforming the Full retrained ACDM in both extrapolation tasks ($Inc_{low}$ and $Inc_{high}$). The deterministic nature of our model handles the predictable vortex shedding highly effectively.
However, in the highly chaotic Transonic Flow regimes (Table 3), the differing strengths of the frameworks become apparent. For the interpolation task ($Tra_{int}$) and near-extrapolation ($Tra_{ext}$), in highly chaotic transonic regimes, diffusion models achieve lower short-term MSE and LSiM by synthesizing fine-scale stochastic features. The diffusion baseline excels here because it can stochastically generate the sharp, high-frequency shock waves characteristic of compressible flows. However, this advantage diminishes over long horizons, where our deterministic, $L_{2}$-optimized KAE exhibits a spectral bias, tending to smooth these chaotic, high-frequency features (frequency-space metrics quantifying this smoothing effect are provided in Appendix E). To better visualize this, difference maps between the ground truth and predictions are included in Appendix F, showing that KAE errors are localized entirely around the shock fronts, whereas ACDM errors are more diffusely distributed.
This smoothing becomes advantageous during the long-horizon rollout ($Tra_{long}$, $T=240$). Over extended autoregressive steps, the stochastic nature of ACDM leads to error accumulation and structural drift, resulting in an MSE of $22.2\times 10^{-3}$. In contrast, the linear latent dynamics remain stable over long horizons. This stability arises from the spectral structure of the learned Koopman operator. As visualized in Appendix E.3, the eigenvalue distribution of the learned operator across several conditioning parameters lies predominantly in the left half of the complex plane, corresponding to dissipative latent dynamics. This structure naturally suppresses unstable growth modes during rollout and helps explain the improved long-horizon stability observed in our experiments. By restricting the latent dynamics to a mathematically stable evolution, our KAE degrades gracefully and achieves a significantly lower MSE of $17.0\times 10^{-3}$, demonstrating superior robustness for extreme long-term forecasting.
For probabilistic models, direct comparison to a single reference trajectory is insufficient, as trajectories naturally decorrelate over time even for accurate simulations (hu2020). To further quantify the temporal degradation of the models, Figure 2 presents the spatial Pearson correlation coefficient between the predicted flow fields and the reference simulations over the 240-step $Tra_{long}$ horizon. While both models maintain high correlation initially, ACDM gradually loses phase alignment with the reference vortex shedding, resulting in a steep drop in correlation and high variance across samples. The Continuous KAE demonstrates a notably slower rate of correlation decay and a tighter variance band, confirming that the linear constraints of the Koopman latent space effectively preserve the underlying temporal phase. A deeper investigation into the spectral properties of this smoothing effect, alongside an extreme 1000-step stress test, is provided in Appendix E.
Figure 2: Spatial Pearson correlation over a 240-step rollout in the Transonic regime ($Tra_{long}$). The Continuous KAE maintains significantly more stable structural alignment compared to the unconstrained diffusion baseline.
@@
We compare trajectories generated using numerical RK4 integration with the analytical matrix exponential solution derived in Section 4.3. While RK4 performs iterative rollouts with a fixed timestep $\Delta t=0.1\,\mathrm{s}$, the analytical formulation allows direct computation of the latent state as $z(\tau)=\exp(K\tau)z_{0}$, for any target time $\tau$. Figure 4 and Table 3 show close agreement between the two approaches, confirming that the learned dynamics are well approximated by the linear generator. This property enables highly efficient long-horizon forecasting, as future states can be computed in a single step without sequential rollout.
@@
A key advantage of our continuous-time KAE is its ability to generalize to arbitrary temporal resolutions without retraining. Despite being trained on a fixed timestep $\Delta t=0.1s$, the model supports arbitrary integration step sizes during inference. We validate this capability in Figure 5, which displays rollouts generated using three distinct integration step sizes, $\Delta t\in\{0.05s,0.1s,0.2s\}$. For each case, the number of rollout steps is adjusted to align with the same set of physical time instants (up to $t=2.0s$). The first row also shows the results produced using the analytical exponential form. The perfect alignment confirms that the learned dynamics are robust to discretization changes.
@@
Architecture Avg. Step (ms) Mean VRAM (MB)
ResNet-m2 $3.67\pm 0.04$ $188.0$
Dil-ResNet-m2 $3.46\pm 0.02$ $\mathbf{178.6}$
FNO-16 $1.17\pm 0.01$ $184.1$
FNO-32 $1.17\pm 0.00$ $183.9$
UNet-m2 $6.19\pm 0.09$ $183.7$
UNet-m8 $6.16\pm 0.01$ $184.1$
TF-Enc $0.60\pm 0.25$ $3448.6$
TF-MGN $0.69\pm 0.01$ $3498.0$
TF-VAE $0.30\pm 0.01$ $13749.9$
Refiner-R4 $10.31\pm 0.02$ $642.4$
KAE $\mathbf{1.04\times 10^{-3}\pm 10^{-4}}$ $2751.3$
ACDM $41.77\pm 0.01$ $659.2$
$\text{ACDM}_{ncn}$ $41.70\pm 0.06$ $649.2$
@@
ResNet $10\pm 9.1$ $17\pm 7.8$ $16\pm 3.0$ $5.9\pm 1.6$ $2.3\pm 0.9$ $1.4\pm 0.2$ $1.8\pm 1.0$ $\mathbf{1.0\pm 0.3}$ $24.2\pm 4.6$ $7.6\pm 1.7$
ResNet-dil $1.6\pm 1.8$ $7.7\pm 5.5$ $2.6\pm 0.7$ $1.7\pm 1.0$ $1.2\pm 0.3$ $1.7\pm 1.4$ $\mathbf{1.0\pm 0.5}$ $5.7\pm 2.1$ $22.0\pm 2.4$ $5.5\pm 2.3$
$\text{FNO}_{16}$ $2.8\pm 3.1$ $8.8\pm 7.1$ $8.9\pm 3.8$ $2.5\pm 1.2$ $4.8\pm 1.2$ $3.4\pm 1.1$ $5.5\pm 2.6$ $2.6\pm 1.1$ $20.8\pm 2.0$ $11.5\pm 1.1$
$\text{FNO}_{32}$ $160\pm 50$ $80\pm 5.4$ $1k\pm 140$ $57\pm 4.9$ $4.9\pm 1.9$ $3.6\pm 0.9$ $6.8\pm 3.4$ $3.1\pm 1.1$ Diverged Diverged
$\text{TF}_{MGN}$ $5.7\pm 4.3$ $13\pm 6.4$ $10\pm 2.9$ $3.5\pm 0.4$ $3.9\pm 1.0$ $1.8\pm 0.3$ $6.3\pm 4.4$ $2.2\pm 0.7$ $18.9\pm 4.5$ $4.0\pm 0.2$
$\text{TF}_{Enc}$ $1.5\pm 1.7$ $6.3\pm 4.2$ $8.7\pm 3.8$ $0.6\pm 0.3$ $\mathbf{1.0\pm 0.3}$ $3.3\pm 1.2$ $1.8\pm 0.3$ $6.2\pm 4.2$ $22.2\pm 3.9$ $3.8\pm 0.4$
$\text{TF}_{VAE}$ $5.4\pm 5.5$ $13\pm 7.2$ $4.1\pm 1.4$ $4.1\pm 0.9$ $2.4\pm 0.2$ $7.2\pm 3.0$ $2.7\pm 0.6$ $11\pm 5.1$ $20.6\pm 2.1$ $4.0\pm 0.2$
U-Net $1.0\pm 1.1$ $5.8\pm 3.2$ $2.7\pm 0.6$ $2.6\pm 0.6$ $3.1\pm 2.1$ $3.9\pm 2.8$ $2.3\pm 2.0$ $3.3\pm 2.8$ $30.3\pm 6.1$ $9.1\pm 1.2$
[…]
$\text{U-Net}_{tn}$ $1.0\pm 1.0$ $5.6\pm 3.1$ $0.9\pm 0.6$ $1.5\pm 0.6$ $4\pm 0.8 1.1$$1\pm 0.3$ $1.8\pm 1.1$ $1.0\pm 0.4$ $22.4\pm 3.9$ $3.9\pm 0.4$
Refiner $1.3\pm 1.4$ $7.1\pm 4.2$ $3.5\pm 2.2$ $2.5\pm 1.0$ $5.4\pm 2.1$ $2.3\pm 0.5$ $7.1\pm 2.1$ $3.0\pm 1.7$ Diverged $8.8\pm 3.3$
$\text{ACDM}_{ncn}$ $0.9\pm 0.8$ $6.6\pm 2.7$ $5.7\pm 2.7$ $3.6\pm 1.2$ $4.1\pm 1.9$ $1.9\pm 0.6$ $2.8\pm 1.3$ $1.7\pm 0.4$ $22.8\pm 3.8$ $8.2\pm 3.0$
ACDM $1.7\pm 2.2$ $6.9\pm 5.6$ $0.8\pm 0.4$ $1.0\pm 0.3$ $2.3\pm 1.4$ $1.3\pm 0.3$ $2.7\pm 2.1$ $1.3\pm 0.6$ $22.6\pm 4.0$ $\mathbf{3.8\pm 0.4}$
Continuous KAE $1.3\pm 1.7$ $6.1\pm 4.8$ $2.9\pm 1.1$ $1.7\pm 0.3$ $2.5\pm 0.8$ $1.8\pm 0.3$ $6.5\pm 1.6$ $2.2\pm 0.4$ $\mathbf{17.0\pm 2.3}$ $4.1\pm 0.3$
Exponential Continuous KAE $1.3\pm 1.7$ $6.1\pm 4.8$ $2.9\pm 1.1$ $1.7\pm 0.3$ $2.5\pm 0.8$ $1.8\pm 0.3$ $6.5\pm 1.6$ $2.2\pm 0.4$ $\mathbf{17.0\pm 2.3}$ $4.1\pm 0.3$
@@
Asymptotic Stability and Dissipation.
To test the absolute limits of the learned operator, we subjected the models to an extreme 1000-step rollout ($T=1000$), shown in Appendix E.2. Over such extended horizons, autoregressive data-space models typically suffer from catastrophic numerical divergence, generating non-physical artifacts. In contrast, our Continuous KAE degrades gracefully. Because the learned latent dynamics are fundamentally governed by a linear Koopman operator, the system is asymptotically stable; the eigenvalues of the operator capture the natural dissipative properties of the fluid. As the rollout progresses towards $t=1000$, the high-frequency transient modes naturally decay, and the prediction smoothly diffuses toward the steady-state base flow. While this results in a loss of fine-scale turbulent texture (quantified by the spectral analysis in Appendix E), it leads to bounded predictions over long horizons, making it highly reliable for extreme long-term structural forecasting. This behavior is consistent with the non-autoregressive nature of our model: because the latent evolution is computed directly via matrix exponentiation, errors do not compound across timesteps.
Figure 6: Visual snapshots of the flow field over an extreme 1000-step rollout in the Transonic regime. While the Continuous KAE smoothly diffuses the flow into a stable limit cycle without blowing up, the unconstrained autoregressive diffusion baseline (ACDM) eventually compounds stochastic errors until the physical structure collapses into numerical noise.
@@ 7 Discussion
In this work, we presented a continuous-time Koopman Autoencoder that effectively decouples the learning of latent dynamics from the temporal resolution of the training data. Our results on incompressible and transonic flows demonstrate that this approach provides an alternative to probabilistic methods like ACDM. They reveal a trade-off between expressivity and dynamical consistency. While generative models capture high-frequency stochastic features, they rely on autoregressive sampling that can lead to instability over long horizons. In contrast, structured latent dynamics impose a strong inductive bias that favors stable and efficient evolution, at the cost of reduced expressivity. This suggests that hybrid approaches combining structured dynamics with expressive residual modeling may offer a promising direction for future work.
Accuracy vs. Efficiency Trade-off. While diffusion models such as ACDM excel at capturing high-frequency texture through stochastic refinement, they incur a significant computational cost during inference due to iterative denoising steps. Our method, by contrast, utilizes a deterministic ODE solver in the latent space. As shown in Table 1, our continuous formulation achieves competitive MSE and LSiM scores, often within the error margins of ACDM, while being two orders of magnitude faster during inference. This efficiency is beneficial for applications requiring fast evaluation.
@@
Spectral Bias and Extreme Long-Horizon Stability. As reflected in our transonic results, the deterministic nature of our continuous-time KAE introduces a spectral bias: it tends to smooth fine-scale, chaotic spatial features. Our frequency analysis (detailed in Appendix E) confirms this, showing a faster energy drop-off at high spatial wavenumbers compared to the diffusion-based ACDM, which synthesizes textural details more accurately. However, the temporal frequency analysis reveals the KAE’s critical advantage: it captures the dominant vortex shedding frequencies with minimal variance. Rather than a strictly negative artifact, this physical low-pass filtering effect is exactly what guarantees the model’s stability. When subjected to an extreme $T=1000$ step rollout, ACDM’s unconstrained generation compounds high-frequency stochastic errors until the physical structure collapses. In contrast, our Continuous KAE remains strictly bounded by its linear latent dynamics, gracefully decaying high-frequency transients to settle into a stable, physically accurate limit cycle without numerical divergence. To address the spectral smoothing while retaining this extreme stability, future hybrid approaches could utilize the continuous-time Koopman state as a mathematically rigorous base flow. A lightweight, conditional diffusion or generative adversarial decoder could then be applied exclusively at the final inference step to stochastically recover fine-scale turbulent textures without risking autoregressive error accumulation.
Limitations. Despite these successes, we observe that the spectral bias of the $L_{2}$ reconstruction loss can lead to the smoothing of fine-scale turbulent features over very long horizons, a common issue in deterministic autoencoders. Additionally, while the Koopman operator provides interpretability, the assumption of global linearity (even with residual corrections) remains a constraint for highly non-linear phenomena like shock wave propagation in the transonic regime.
@@ 8 Conclusion
We introduced a continuous-time Koopman Autoencoder for fluid dynamics forecasting, modelling latent evolution as an ODE governed by a parameter-conditioned linear generator. This formulation combines the interpretability of Koopman theory with the flexibility of neural differential equations, enabling accurate, zero-shot temporal-resolution-invariant predictions. Extensive benchmarks on incompressible and transonic flows confirm that our approach offers massive computational efficiency and extreme long-horizon stability compared to highly expressive autoregressive diffusion models.
Future work will scale this approach to three-dimensional turbulence—made computationally viable by our low-rank parameter adaptation—and incorporate explicit conservation laws into the latent ODE to further improve physical fidelity. Ultimately, this work demonstrates that enforcing structured, continuous-time linear dynamics in latent spaces provides a scalable, efficient, and robust alternative for complex PDE forecasting where long-term stability is critical.
This work demonstrates that enforcing simple linear structure in latent space can yield robust and scalable models for dynamical systems. By combining Koopman theory with continuous-time dynamics, we obtain models that are temporally consistent, computationally efficient, and stable over long horizons. These results suggest that structured latent representations provide a viable alternative to highly expressive autoregressive models, particularly in regimes where stability and generalization are critical. More broadly, this work highlights the value of combining physical structure with learned representations to build scalable and reliable models for complex dynamical systems.
@@
This paper introduces a surrogate model that significantly accelerates fluid dynamics simulations. By reducing the reliance on computationally expensive numerical solvers (DNS/LES), this method facilitates more energy-efficient workflows in engineering and climate science. This contributes to lowering the carbon footprint associated with high-performance computing, with no foreseeable negative ethical or societal risks.
@@
Datasets are publicly available at https://github.com/tum-pbs/autoreg-pde-diffusion/tree/main at the training and evaluation resolution of $128\times 64$ ($\sim$146 GB).